A rock is dropped from a height of 2.7 m. How fast is it going when it hits the ground?


1
Expert's answer
2021-05-18T11:10:10-0400

We can find the velocity of the rock when it hits the ground from the law of conservation of energy:


"PE=KE,""mgh=\dfrac{1}{2}mv^2,""v=\sqrt{2gh}=\sqrt{2\cdot9.8\ \dfrac{m}{s^2}\cdot2.7\ m}=7.3\ \dfrac{m}{s}."